### v1

@@ Appendix B Training Implementation Details
This section provides implementation details for achieving the theoretical guarantees established in our analysis, focusing on the demand predictor design and constrained training framework that ensures the performance advantage condition is satisfied.
B.1 Demand Predictor Implementation
The demand predictor $F_{t}=\text{Predictor}(U_{t-K:t},Q_{t-K:t},\\ H_{t-K:t})$ employs a simple yet effective multi-layer perceptron architecture. The input consists of concatenated historical features over $K=5$ timeslots, yielding $\text{Input}=\text{Concat}([U_{t-5:t},Q_{t-5:t},H_{t-5:t}])\in\mathbb{R}^{5\times 3R}$. The network architecture follows a standard 3-layer design with ReLU activations: input layer ($15R$ dimensions) $\rightarrow$ hidden layer (512 units) $\rightarrow$ hidden layer (256 units) $\rightarrow$ output layer ($R$ dimensions with softmax). The training objective minimizes mean squared error with L2 regularization: $\mathcal{L}_{\text{pred}}=\frac{1}{N}\sum_{i=1}^{N}||F_{t+1}^{\text{pred}}-F_{t+1}^{\text{actual}}||_{2}^{2}+\lambda||\theta||_{2}^{2}$ where $\lambda=10^{-4}$.
B.2 Constrained Training Framework
To ensure the performance advantage condition $\frac{1-1/s}{\epsilon}>\frac{L_{R}+\beta L_{P}}{\alpha\cdot K_{0}}$ is satisfied during training, we implement adaptive constraint enforcement. The constraint weights $\gamma_{t}$ and $\delta_{t}$ for terms $\mathcal{L}_{\epsilon}$ and $\mathcal{L}_{s}$ are dynamically adjusted based on current performance: $\gamma_{t}=\gamma_{0}\cdot\exp(\alpha_{\gamma}\cdot\max(0,||B_{t}||_{F}-\epsilon_{\text{target}}))$ and $\delta_{t}=\delta_{0}\cdot\exp(\alpha_{\delta}\cdot\max(0,s_{\text{target}}-s_{\text{current}}))$, where $B_{t}=A_{t}^{\text{RL}}-A_{t}^{\text{OT}}$ represents the deviation from the optimal transport solution.
@@ B.2 Constrained Training Framework
/* Initialization and Baseline Parameter Estimation */
@@ 3 Pre-train predictor $P_{\psi}$ on historical data $\mathcal{D}$;
/* Estimate Baseline Parameters for Theoretical Conditions */
@@ 6 $L_{R},L_{P}\leftarrow$ Estimate Lipschitz constants via finite differences;
       /* Collect Experience with Constraint Monitoring */
@@ 15             Update constraint weights $\gamma_{t},\delta_{t}$ based on current violations;
      /* Policy Update with Theoretical Constraints */
@@ 18       Compute PPO loss $\mathcal{L}_{\text{PPO}}$ from collected experience;
19       Compute constraint losses $\mathcal{L}_{\epsilon}=\max(0,\frac{||B_{t}||_{F}-\epsilon_{\text{target}}}{\epsilon_{0}})$;
@@
      /* Gradient Update and Condition Validation */

### v2

@@ Appendix B Training Implementation Details
// Initialization and Baseline Parameter Estimation
@@ 3 Pre-train predictor $P_{\psi}$ on historical data $\mathcal{D}$;
// Estimate Baseline Parameters for Theoretical Conditions
@@ 6 $L_{R},L_{P}\leftarrow$ Estimate Lipschitz constants via finite differences;
  // Collect Experience with Constraint Monitoring
@@ 15     Update constraint weights $\gamma_{t},\delta_{t}$ based on current violations;
 // Policy Update with Theoretical Constraints
@@ 18  Compute PPO loss $\mathcal{L}_{\text{PPO}}$ from collected experience;
[…]
@@
 // Gradient Update and Condition Validation
@@
B-A Network Details
For Reinforcement learning, the policy network $\pi_{\theta}$ is implemented as a fully connected neural network with three hidden layers (256, 512, 256 units) and ReLU activations. The output layer uses a softmax activation to generate probability distributions over regional allocations. The value network $V_{\phi}$ shares the same architecture but outputs scalar state values. We use the Adam optimizer with an initial learning rate of $3\times 10^{-4}$ and exponential decay with factor 0.995 every 100 episodes.
[…]
B-B Constrained Training Framework
[…]